OpenStudy (compassionate):

Solve this and receive a free Pepsi Max. 48÷ 2(9 + 3) = ?

hartnn (hartnn):

its 288 because we go from left to right if there's only * and / (ofcourse doing 9+3=12 first) though calcy gives 2.

OpenStudy (anonymous):

Use the order of operations. 48÷ 2(9 + 3) 48÷ 2(12) 48÷24 2

OpenStudy (anonymous):

Now hopefully, there is no ambiguity here and it isn't asking for 48 ÷ 2 × (9+3), then the division would have to be done first . . .

OpenStudy (compassionate):

24(9 + 3) =?

OpenStudy (anonymous):

NO! I am correct! Trust me! I am a math major!

OpenStudy (compassionate):

I got my major 20 years ago.

hartnn (hartnn):

there would be no ambiguity if it was written like this:48÷ (2(9 + 3))

OpenStudy (anonymous):

Why did you ask the question then?

OpenStudy (compassionate):

It's a troll question that is rather fun to watch people debate.

OpenStudy (zzr0ck3r):

silly

hartnn (hartnn):

thats the reason i went with 288 if everyone says 2, whom should i debate wit ?? :P

hartnn (hartnn):

here's the proof: http://www.wolframalpha.com/input/?i=48%C3%B72%289+%2B+3%29

OpenStudy (anonymous):

Troll on, troller. \m/

hartnn (hartnn):

exactly 'interpretation of notation' i can say dillpickles incorrectly interpreted it as \(\frac{48}{2(9+3)}\) and i correctly interpreted it as \(\frac{48(9+3)}{2}\)

OpenStudy (anonymous):

For the record, my calculator gives 288 as the result, but it's just a machine and doesn't know any better.

hartnn (hartnn):

about human made rules,what i have heard is to use PEDMAS and when both D and M comes,u go from left to right,that gives 288.

OpenStudy (anonymous):

I'm sure I've said this here before, but Moral of the story: When in doubt, use more parentheses.

hartnn (hartnn):

^^ thats the same thing i told: 'there would be no ambiguity if it was written like this:48÷ (2(9 + 3))'

OpenStudy (anonymous):

Yep. :-) With PEMDAS, I make a distinction between 2(9+3) and 2 * (9+3). If the product 2(9+3) is a separate term like the statement was originally presented, then I treat it as a separate term. 2 * (9+3) makes it more clear that the operations are to be done in order. Mathematics is about learning new facts in order to answer questions, sometimes the best answer to give is: "The question was not properly posed."